# Static and Dynamic BART for Rank-Order Data

### v4

@@ 3.1 Prior on the BART Parameters
where $\mu_{\ell,s}\in\bm{\mu}_{s}$. To deal with the potential overfitting of the nonparametric BART specification, we follow Chipman et al. (2010) and choose a regularization prior for the tree structure and terminal nodes that constrains the size and fit of each tree so that each one contributes only a small part to the overall fit. This prior strongly encourages each tree to be a weak learner, with leaf parameters close to zero.
The marginal prior for each tree structure, $p(\mathcal{T}_{s})$, is specified by three parts. First, the probability that a node at depth $d=0,1,\ldots$ is nonterminal, given by $\underline{\alpha}(1+d)^{-\underline{\beta}}$, for $\underline{\alpha}\in(0,1)$, $\underline{\beta}\in[0,\infty)$. Smaller (larger) values of $\underline{\alpha}$ ($\underline{\beta}$) impose a larger penalty on more complex tree structures. We use the default choice proposed by Chipman et al. (2010), that is $\underline{\alpha}=0.95$ and $\underline{\beta}=2$. Second, the distribution of the splitting variable at each interior node is assumed to be discrete uniform over all the covariates. Third, the distribution on the splitting rule assignment in each interior node, $c_{h}$, conditional on the splitting variable, is uniform on the discrete set of available splitting values.
The conditional prior on the terminal leaf parameters, $p(\mu_{\ell,s}|\mathcal{T}_{s})$, is assumed to be the conjugate normal distribution $\mathcal{N}(\mu_{\mu},\sigma^{2}_{\mu})$, where the hyperparameters are chosen such that $\mathcal{N}(S\mu_{\mu},S\sigma^{2}_{\mu})$ assigns substantial probability to the interval $(z_{min},z_{max})$. In line with Chipman et al. (2010), the prior on $\mu_{\ell,s}$ is constructed with $\mu_{\mu}=0$ and $\sigma_{\mu}=(z_{max}-z_{min})/(2k_{\mu}\sqrt{S})$, where $k_{\mu}$ is a hyperparameter set to 2, and $z_{max}$ and $z_{min}$ are the maximum and minimum of initial values of the latent outcome.
@@ 3.2 Posterior Sampling for ARROBART
    sample the path of the latent variables $\mathbf{z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{T}\}$, recursively over time;
@@
    sample the tree and leaf parameters from $p(\mathcal{T}_{s},\bm{\mu}_{s}|\mathcal{T}_{-s},\bm{\mu}_{-s},\mathbf{z})$, for $s=1,\ldots,S$:
@@
Sampling the path of the latent variables using the filtering and smoothing recursions in Theorem 1 is highly computationally intensive. This is due to (i) evaluation of many Gaussian integrals over very small subspaces of $\mathds{R}^{N}$ and (ii) drawing from an $N$-dimensional Gaussian distribution truncated such that its elements satisfy a given ordering constraint. Both issues worsen as the number of items, $N$, increases. To overcome these computational issues, for each period $t=1,\ldots,T$, we approximate the joint posterior distribution of the vector $\mathbf{z}_{t}$ and derive the element-wise posterior full conditional distribution for each $i=1,\ldots,N$, conditioning on all the other $j\neq i$ elements. Therefore, we replace sampling from a joint distribution by sampling from a collection of full conditional distributions, which results in a full Gibbs sampler of the latent scores. This approach is similar to the standard technique used in rank-order data models (e.g., Hoff, 2009, Ch. 12). We refer to the Supplement for a detailed description of the sampling method.
Steps 2b and 2c concerning the static parameters of the tree structure are standard in the BART literature and follow directly from the algorithm in Chipman et al. (2010). The tree structure is updated by proposing one of four moves: growing a terminal node, pruning a pair of terminal nodes, changing a nonterminal rule, and swapping a rule between parent and child. Then, given the tree structure, the terminal node parameters are sampled directly from their normal conditional posterior distribution.
@@ 4 Simulation Studies
This section investigates the performance of the proposed methods on synthetic data. In particular, Section 4.1 compares the ROBART model to the methods for rank-order data introduced by Li et al. (2022). Then, Section 4.2 presents a comparative study between ARROBART and a range of models that account for (or ignore) the possible temporal persistence of the data.
@@ 4.1.1 Static Simulation Study with Full Rankings
The performance of the ROBART model defined in Section 2.2 is compared to the methods for rank-order data introduced by Li et al. (2022) by using the normalized Kendall tau distance (Kendall, 1938). Given two ranking lists $\hat{\tau}$ and $\tau$, the normalized Kendall tau distance is defined as the proportion of pairwise disagreements between them:
@@
The three synthetic datasets used in this simulation are generated as in Li et al. (2022), and we report briefly their structures. For each $i=1,\ldots,N$, item $i$ has a true score $\gamma_{i}$, while the covariate vectors $\mathbf{x}_{i}=(x_{i1},\ldots,x_{iK_{x}})^{\prime}$ are drawn independently from a Normal distribution with zero mean and covariance equal to $\mathbb{C}ov(x_{il},x_{im})=\rho^{\left|l-m\right|}$, for $1\leq l,m\leq v$ and $\left|\rho\right|<1$. The true score vector $\bm{\gamma}$ is generated by considering the different roles of the covariates as:
@@
For each scenario and each value of $\sigma\in\{1,5,10,20,40\}$, we generate $100$ synthetic datasets with $N=50$ and $M=10$. We compare the results for ROBART against several competitors: the Borda Count, the Markov-Chain methods (MC1, MC2, MC3 - Dwork et al., 2001), the Cross-Entropy Monte Carlo method (CEMC - Lin and Ding, 2009), the Plankett-Luce (PL) model, the BARC model with and without covariates, and the BARC model with Mixture of rankers with different opinions (BARCM - Li et al., 2022). Further details on these methods are available in the Supplement.
We use the default values for the ROBART model as in Chipman et al. (2010) and follow Li et al. (2022) in choosing the hyperparameters for all the remaining models.
$\sigma$ Borda MC1 MC2 MC3 CEMC PL BAR $\text{BARC}_{1}$ $\text{BARC}_{2}$ $\text{BARC}_{3}$ BARCM ROBART Scenario 1 1 (0.03) 1.31 1.08 1.01 2.55 5.62 0.99 0.65 0.65 0.98 0.65 0.89 5 (0.13) 1.27 1.02 1.01 1.13 1.09 0.99 0.65 0.65 0.97 0.65 0.68 10 (0.22) 1.57 1.01 1.00 1.05 1.07 1.00 0.71 0.71 0.99 0.71 0.69 20 (0.32) 1.47 1.00 1.00 1.02 1.04 0.99 0.78 0.78 0.99 0.78 0.75 40 (0.41) 1.23 1.00 1.00 1.01 1.01 0.99 0.88 0.88 0.99 0.88 0.87 Scenario 2 1 (0.03) 1.33 1.04 1.00 2.65 1.06 0.99 0.97 0.97 0.98 0.97 0.86 5 (0.12) 1.30 1.01 1.00 1.12 1.10 0.99 0.83 0.83 0.97 0.83 0.69 10 (0.20) 1.36 1.01 1.00 1.05 1.07 1.00 0.77 0.77 0.99 0.77 0.67 20 (0.29) 1.48 1.01 1.00 1.03 1.04 0.99 0.80 0.80 0.99 0.80 0.72 40 (0.39) 1.25 1.00 1.00 1.01 1.03 1.00 0.89 0.89 1.00 0.89 0.85 Scenario 3 1 (0.05) 1.28 1.03 1.00 1.60 1.11 1.00 1.00 1.00 1.00 1.00 0.88 5 (0.18) 1.32 1.01 1.00 1.06 1.06 0.99 1.00 1.00 0.99 1.00 0.73 10 (0.29) 1.38 1.00 1.00 1.03 1.04 0.99 1.00 1.00 0.99 1.00 0.78 20 (0.36) 1.34 1.00 1.00 1.02 1.03 1.00 1.00 1.00 1.00 1.00 0.85 40 (0.42) 1.18 1.00 1.00 1.01 1.01 1.00 1.00 1.00 1.00 1.00 0.91  Notes: Scaled Kendall tau distance of rank predictions for the scenarios specified by Li et al. (2022). The Borda column with parentheses shows the average Kendall tau distances between estimated and true ranking lists for the Borda Count predictions, while the remaining columns show tau distances relative to the tau distance attained by the Borda Count method. Bold denotes the best-performing model.
@@ 4.1.2 Static Simulation Study with Partial Rankings
ROBART and ARROBART are extendable to partial ranking data by replacing the constraints on the latent scores with the less restrictive constraints implied by partial rankings. The Supplement contains additional details for the model and sampler for partial ranking data. We demonstrate the ability of our model to predict rankings using partial ranking information through a simulation study described by Li et al. (2022). The comparison also included two additional specifications: the RODART, which is ROBART with a Dirichlet hyperprior on covariates’ splitting probabilities, and the SoftROBART, which is ROBART with soft and sparse splitting rules (Linero and Yang, 2018).
We divide the $N=80$ items into $K=1,2,4,8,10,16$ groups, of $N/K$ items. Pairwise comparison information is only available within groups. The DGP is otherwise the same as scenario 2 above. Table 2 contains the correlations of rank predictions averaged over 10 repetitions of the simulation study with latent error standard deviations $\sigma\in\{1,5\}$, as in Li et al. (2022). The results confirm that ROBART still provides a considerable improvement over linear models when only partial rankings are available. Finally, the SoftROBART is found to provide further improvements in ranking predictions.
K BAR BARC1 BARC2 BARC3 ROBART RODART SoftROBART $\sigma=1$ Scenario 2 1 (0.03) 1.00 1.00 1.00 0.83 0.84 0.67 2 (0.07) 0.61 0.60 0.59 0.32 0.34 0.28 4 (0.08) 0.59 0.62 0.64 0.35 0.39 0.28 8 (0.13) 0.51 0.49 0.56 0.24 0.27 0.21 10 (0.12) 0.59 0.59 0.65 0.26 0.33 0.22 16 (0.17) 0.55 0.56 0.58 0.23 0.36 0.21 $\sigma=5$ Scenario 2 1 (0.12) 0.83 0.83 0.98 0.60 0.68 0.46 2 (0.13) 0.80 0.79 1.04 0.54 0.64 0.46 4 (0.14) 0.78 0.78 0.95 0.56 0.73 0.46 8 (0.17) 0.63 0.64 0.84 0.46 0.58 0.36 10 (0.16) 0.65 0.65 0.88 0.45 0.57 0.37 16 (0.19) 0.60 0.60 0.78 0.42 0.58 0.35  Notes: Scaled Kendall tau distance of rank predictions for the scenarios specified by Li et al. (2022). The BAR column with parentheses shows the average Kendall tau distances between estimated and true ranking lists for the BAR predictions, while the remaining columns show tau distances relative to the tau distance attained by BAR.
@@ 4.1.3 Static Simulation Study with Ranker-specific Covariates
Table  3 contains the results for a variation on the static simulation study described by Li et al. (2022) in which the covariates are ranker-specific. For each $(i,m)\in\{1,\ldots,N\}\times\{1,\dots,M\}$, item $(i,m)$ has a true score $\gamma_{i,m}$, while the covariate vectors $\mathbf{x}_{i,m}=(x_{i,m,1},\ldots,x_{i,m,K_{x}})^{\prime}$ are drawn independently from a Normal distribution with zero mean and covariance equal to $\mathbb{C}ov(x_{i,m,l},x_{i,m,s})=\rho^{\left|l-s\right|}$, for $1\leq l,s\leq v$ and $\left|\rho\right|<1$. The true score vector $\bm{\gamma}$ is generated by considering the different roles of the covariates as:
@@ 4.2 Comparison of ARROBART and Linear Models
where in the last scenario the $K_{x}$ covariates $\mathbf{x}_{ij,t}=(x_{ij,t,1},\ldots,x_{ij,t,K_{x}})^{\prime}$ are drawn independently from a multivariate normal distribution with mean zero and $\mathbb{C}ov(x_{ij,t,l},x_{ij,t,m})=\rho^{\left|l-m\right|}$, for $1\leq l,m\leq v$ and $\left|\rho\right|<1$. In detail, each scenario describes a nonlinear relationship, where $i=1,\ldots,N$, $j=1,\ldots,M$, and $t=1,\ldots,T$. We fix the number of rankers, $M$, the number of items, $N$ and the period, $T$, to $5$, $20$ and $52$, respectively.
@@
We compare the results for the full and the partial ranking against different competitors, such as (i) an autoregressive linear model for the latent score (ARROLinear); (ii) an ARROBART specification with exogenous covariates (ARROBARTX); (iii) an autoregressive linear model with exogenous covariates (ARROLinearX). Moreover, we extend the ARROBART model to include the lag-1 of the observed rank (ARROBART-lag) and compare it with (i) an analogous linear model (ARROLinear-lag); (ii) a ROBART model using the first lag of the response variable as a covariate (ROBART-lag); (iii) a linear rank order model (BARC) with the first lag of the response variable (ROLinear-lag).
@@
The results are shown in Table 4, where the normalized Kendall tau distance is reported for the benchmark ARROBART model, whereas for the competitors the ratio of their performance over the ARROBART is shown. Instead, Table 5 illustrates the outcome for all the models estimated using also the lag of the observed rank as a covariate, where we use the ARROBART-lag model as a benchmark. For the benchmark models, the closer the normalized Kendall tau distance is to zero, the better the performance. Therefore, ratios of Kendall’s tau distances for the competitors greater than 1 mean that the benchmark is outperforming them (and vice versa). The results suggest several interesting insights. First, the ARROBART (ARROBART-lag) benchmark outperforms the competitors for almost all the noise levels in scenarios 1 and 2. In scenarios 1 and 2, ARROBART-lag is almost uniformly better than all methods. The improvement over ARROLinear and Kendall’s tau for ARROBART are worse in the presence of a strong signal ($\sigma=0.1$); this is possibly due to the sub-optimal performance of the BART specification in almost the absence of noise and is in line with previous results in the literature (Chipman et al., 2010). Second, in the third scenario, the ARROBARTX (Tab. 4) and ROBART-lag (Tab. 5) are the best models. This demonstrates the ability of the proposed methods to leverage exogenous covariates when they are relevant to the regression. Third, the results suggest that the nonlinear specifications, that is the ARROBART and ROBART models, almost always outperform their respective linear counterparts. Finally, the results from Tab. 4 and  5 are confirmed even in the presence of partial ranking.

### v5

@@ 3.1 Prior on the BART Parameters
where $\mu_{\ell,s}\in\bm{\mu}_{s}$. To deal with the potential overfitting of the nonparametric BART specification, we follow chipman2010bart and choose a regularization prior for the tree structure and terminal nodes that constrains the size and fit of each tree so that each one contributes only a small part to the overall fit. This prior strongly encourages each tree to be a weak learner, with leaf parameters close to zero.
The marginal prior for each tree structure, $p(\mathcal{T}_{s})$, is specified by three parts. First, the probability that a node at depth $d=0,1,\ldots$ is nonterminal, given by $\underline{\alpha}(1+d)^{-\underline{\beta}}$, for $\underline{\alpha}\in(0,1)$, $\underline{\beta}\in[0,\infty)$. Smaller (larger) values of $\underline{\alpha}$ ($\underline{\beta}$) impose a larger penalty on more complex tree structures. We use the default choice proposed by chipman2010bart, that is $\underline{\alpha}=0.95$ and $\underline{\beta}=2$. Second, the distribution of the splitting variable at each interior node is assumed to be discrete uniform over all the covariates. Third, the distribution on the splitting rule assignment in each interior node, $c_{h}$, conditional on the splitting variable, is uniform on the discrete set of available splitting values.
The conditional prior on the terminal leaf parameters, $p(\mu_{\ell,s}|\mathcal{T}_{s})$, is assumed to be the conjugate normal distribution $\mathcal{N}(\mu_{\mu},\sigma^{2}_{\mu})$, where the hyperparameters are chosen such that $\mathcal{N}(S\mu_{\mu},S\sigma^{2}_{\mu})$ assigns substantial probability to the interval $(z_{min},z_{max})$. In line with chipman2010bart, the prior on $\mu_{\ell,s}$ is constructed with $\mu_{\mu}=0$ and $\sigma_{\mu}=(z_{max}-z_{min})/(2k_{\mu}\sqrt{S})$, where $k_{\mu}$ is a hyperparameter set to 2, and $z_{max}$ and $z_{min}$ are the maximum and minimum of initial values of the latent outcome.
@@ 3.2 Posterior Sampling for ARROBART
    given the static parameters, sample the latent states $\mathbf{z}=\{\mathbf{z}_{1},\ldots,\mathbf{z}_{T}\}$;
@@
    given the latent states, sample the tree and leaf parameters from $p(\mathcal{T}_{s},\bm{\mu}_{s}|\mathcal{T}_{-s},\bm{\mu}_{-s},\mathbf{z})$, for $s=1,\ldots,S$:
@@
The latent states can be sampled through forward filtering backward sampling, using the filtering recursions in Theorem 1 and a backward sampler described in the Supplement. However, this sampler is computationally intensive. This is due to (i) evaluation of many Gaussian integrals over very small subspaces of $\mathds{R}^{N}$ and (ii) drawing from an $N$-dimensional Gaussian distribution truncated such that its elements satisfy a given ordering constraint. Both issues worsen as the number of items, $N$, increases. To overcome these computational issues, for each period $t=1,\ldots,T$, we approximate the joint posterior distribution of the vector $\mathbf{z}_{t}$ and derive the element-wise posterior full conditional distribution for each $i=1,\ldots,N$, conditioning on all the other $j\neq i$ elements. Therefore, we replace sampling from a joint distribution by sampling from a collection of full conditional distributions, which results in a full Gibbs sampler of the latent scores. This approach is similar to the standard technique used in rank-order data models (e.g., hoff2009first, Ch. 12). We refer to the Supplement for a detailed description of the sampling method.
Steps 2b and 2c concerning the static parameters of the tree structure are standard in the BART literature and follow directly from the algorithm in chipman2010bart. The tree structure is updated by proposing one of four moves: growing a terminal node, pruning a pair of terminal nodes, changing a nonterminal rule, and swapping a rule between parent and child. Then, given the tree structure, the terminal node parameters are sampled directly from their normal conditional posterior distribution.
@@ 4 Simulation Studies
This section investigates the performance of the proposed methods on synthetic data. In particular, Section 4.1 compares the ROBART model to the methods for rank-order data introduced by li2016bayesian. Then, Section 4.2 presents a comparative study between ARROBART and a range of models that account for (or ignore) the possible temporal persistence of the data.
@@ 4.1.1 Static Simulation Study with Full Rankings
The performance of the ROBART model defined in Section 2.2 is compared to the methods for rank-order data introduced by li2016bayesian by using the normalized Kendall tau distance (kendall1938new). Given two ranking lists $\hat{\tau}$ and $\tau$, the normalized Kendall tau distance is defined as the proportion of pairwise disagreements between them:
@@
The three synthetic datasets used in this simulation are generated as in li2016bayesian, and we report briefly their structures. For each $i=1,\ldots,N$, item $i$ has a true score $\gamma_{i}$, while the covariate vectors $\mathbf{x}_{i}=(x_{i1},\ldots,x_{iK_{x}})^{\prime}$ are drawn independently from a Normal distribution with zero mean and covariance equal to $\mathbb{C}ov(x_{il},x_{im})=\rho^{\left|l-m\right|}$, for $1\leq l,m\leq v$ and $\left|\rho\right|<1$. The true score vector $\bm{\gamma}$ is generated by considering the different roles of the covariates as:
@@
For each scenario and each value of $\sigma\in\{1,5,10,20,40\}$, we generate $100$ synthetic datasets with $N=50$ and $M=10$. We compare the results for ROBART against several competitors: the Borda Count, the Markov-Chain methods (MC1, MC2, MC3 - dwork2001rank), the Cross-Entropy Monte Carlo method (CEMC - lin2009integration), the Plankett-Luce (PL) model, the BARC model with and without covariates, and the BARC model with Mixture of rankers with different opinions (BARCM - li2016bayesian). Further details on these methods are available in the Supplement.
We use the default values for the ROBART model as in chipman2010bart and follow li2016bayesian in choosing the hyperparameters for all the remaining models.
$\sigma$ Borda MC1 MC2 MC3 CEMC PL BAR $\text{BARC}_{1}$ $\text{BARC}_{2}$ $\text{BARC}_{3}$ BARCM ROBART Scenario 1 1 (0.03) 1.31 1.08 1.01 2.55 5.62 0.99 0.65 0.65 0.98 0.65 0.89 5 (0.13) 1.27 1.02 1.01 1.13 1.09 0.99 0.65 0.65 0.97 0.65 0.68 10 (0.22) 1.57 1.01 1.00 1.05 1.07 1.00 0.71 0.71 0.99 0.71 0.69 20 (0.32) 1.47 1.00 1.00 1.02 1.04 0.99 0.78 0.78 0.99 0.78 0.75 40 (0.41) 1.23 1.00 1.00 1.01 1.01 0.99 0.88 0.88 0.99 0.88 0.87 Scenario 2 1 (0.03) 1.33 1.04 1.00 2.65 1.06 0.99 0.97 0.97 0.98 0.97 0.86 5 (0.12) 1.30 1.01 1.00 1.12 1.10 0.99 0.83 0.83 0.97 0.83 0.69 10 (0.20) 1.36 1.01 1.00 1.05 1.07 1.00 0.77 0.77 0.99 0.77 0.67 20 (0.29) 1.48 1.01 1.00 1.03 1.04 0.99 0.80 0.80 0.99 0.80 0.72 40 (0.39) 1.25 1.00 1.00 1.01 1.03 1.00 0.89 0.89 1.00 0.89 0.85 Scenario 3 1 (0.05) 1.28 1.03 1.00 1.60 1.11 1.00 1.00 1.00 1.00 1.00 0.88 5 (0.18) 1.32 1.01 1.00 1.06 1.06 0.99 1.00 1.00 0.99 1.00 0.73 10 (0.29) 1.38 1.00 1.00 1.03 1.04 0.99 1.00 1.00 0.99 1.00 0.78 20 (0.36) 1.34 1.00 1.00 1.02 1.03 1.00 1.00 1.00 1.00 1.00 0.85 40 (0.42) 1.18 1.00 1.00 1.01 1.01 1.00 1.00 1.00 1.00 1.00 0.91 Notes: Scaled Kendall tau distance of rank predictions for the scenarios specified by li2016bayesian. The Borda column with parentheses shows the average Kendall tau distances between estimated and true ranking lists for the Borda Count predictions, while the remaining columns show tau distances relative to the tau distance attained by the Borda Count method. Bold denotes the best-performing model.
@@ 4.1.2 Static Simulation Study with Partial Rankings
ROBART and ARROBART are extendable to partial ranking data by replacing the constraints on the latent scores with the less restrictive constraints implied by partial rankings. The Supplement contains additional details for the model and sampler for partial ranking data. We demonstrate the ability of our model to predict rankings using partial ranking information through a simulation study described by li2016bayesian. The comparison also included two additional specifications: the RODART, which is ROBART with a Dirichlet hyperprior on covariates’ splitting probabilities, and the SoftROBART, which is ROBART with soft and sparse splitting rules (linero2018dart).
We divide the $N=80$ items into $K=1,2,4,8,10,16$ groups, of $N/K$ items. Pairwise comparison information is only available within groups. The DGP is otherwise the same as scenario 2 above. Table S.1 contains the correlations of rank predictions averaged over 10 repetitions of the simulation study with latent error standard deviations $\sigma\in\{1,5\}$, as in li2016bayesian. The results confirm that ROBART still provides a considerable improvement over linear models when only partial rankings are available. Finally, the SoftROBART is found to provide further improvements in ranking predictions.
K BAR BARC1 BARC2 BARC3 ROBART RODART SoftROBART $\sigma=1$ Scenario 2 1 (0.03) 1.00 1.00 1.00 0.83 0.84 0.67 2 (0.07) 0.61 0.60 0.59 0.32 0.34 0.28 4 (0.08) 0.59 0.62 0.64 0.35 0.39 0.28 8 (0.13) 0.51 0.49 0.56 0.24 0.27 0.21 10 (0.12) 0.59 0.59 0.65 0.26 0.33 0.22 16 (0.17) 0.55 0.56 0.58 0.23 0.36 0.21 $\sigma=5$ Scenario 2 1 (0.12) 0.83 0.83 0.98 0.60 0.68 0.46 2 (0.13) 0.80 0.79 1.04 0.54 0.64 0.46 4 (0.14) 0.78 0.78 0.95 0.56 0.73 0.46 8 (0.17) 0.63 0.64 0.84 0.46 0.58 0.36 10 (0.16) 0.65 0.65 0.88 0.45 0.57 0.37 16 (0.19) 0.60 0.60 0.78 0.42 0.58 0.35 Notes: Scaled Kendall tau distance of rank predictions for the scenarios specified by li2016bayesian. The BAR column with parentheses shows the average Kendall tau distances between estimated and true ranking lists for the BAR predictions, while the remaining columns show tau distances relative to the tau distance attained by BAR.
@@ 4.1.3 Static Simulation Study with Ranker-specific Covariates
Table S.2 contains the results for a variation on the static simulation study described by li2016bayesian in which the covariates are ranker-specific. For each $(i,m)\in\{1,\ldots,N\}\times\{1,\dots,M\}$, item $(i,m)$ has a true score $\gamma_{i,m}$, while the covariate vectors $\mathbf{x}_{i,m}=(x_{i,m,1},\ldots,x_{i,m,K_{x}})^{\prime}$ are drawn independently from a Normal distribution with zero mean and covariance equal to $\mathbb{C}ov(x_{i,m,l},x_{i,m,s})=\rho^{\left|l-s\right|}$, for $1\leq l,s\leq v$ and $\left|\rho\right|<1$. The true score vector $\bm{\gamma}$ is generated by considering the different roles of the covariates as:
@@ 4.2 Comparison of ARROBART and Linear Models
where in the last scenario the $K_{x}$ covariates $\mathbf{x}_{ij,t}=(x_{ij,t,1},\ldots,x_{ij,t,K_{x}})^{\prime}$ are drawn independently from a multivariate normal distribution with mean zero and $\mathbb{C}ov(x_{ij,t,l},x_{ij,t,m})=\rho^{\left|l-m\right|}$, for $1\leq l,m\leq v$ and $\left|\rho\right|<1$. In detail, each scenario describes a nonlinear relationship, where $i=1,\ldots,N$, $j=1,\ldots,M$, and $t=1,\ldots,T$. We fix the number of rankers, $M$, the number of items, $N$ and the period, $T$, to $5$, 2020 and $52$, respectively.
@@
We compare the results for the full and the partial ranking against different competitors, such as (i) an autoregressive linear model for the latent score (ARROLinear); (ii) an ARROBART specification with exogenous covariates (ARROBARTX); (iii) an autoregressive linear model with exogenous covariates (ARROLinearX). Moreover, we extend the ARROBART model to include the lag-1 of the observed rank (ARROBART-lag) and compare it with (i) an analogous linear model (ARROLinear-lag); (ii) a ROBART model using the first lag of the observed rank as a covariate (ROBART-lag); (iii) a linear rank order model (BARC) with the first lag of the observed rank (ROLinear-lag).
@@
The results are shown in Table 4, where the normalized Kendall tau distance is reported for the benchmark ARROBART model, whereas for the competitors the ratio of their performance over the ARROBART is shown. Instead, Table 5 illustrates the outcome for all the models estimated using also the lag of the observed rank as a covariate, where we use the ARROBART-lag model as a benchmark. For the benchmark models, the closer the normalized Kendall tau distance is to zero, the better the performance. Therefore, ratios of Kendall’s tau distances for the competitors greater than 1 mean that the benchmark is outperforming them (and vice versa). The results suggest several interesting insights. First, the ARROBART (ARROBART-lag) benchmark outperforms the competitors for almost all the noise levels in scenarios 1 and 2. In scenarios 1 and 2, ARROBART-lag is almost uniformly better than all methods. The improvement over ARROLinear and Kendall’s tau for ARROBART are worse in the presence of a strong signal ($\sigma=0.1$); this is possibly due to the sub-optimal performance of the BART specification in almost the absence of noise and is in line with previous results in the literature (chipman2010bart). Second, in the third scenario, the ARROBARTX (Tab. 4) and ROBART-lag (Tab. 5) are the best models. This demonstrates the ability of the proposed methods to leverage exogenous covariates when they are relevant to the regression. Third, the results suggest that the nonlinear specifications, that is the ARROBART and ROBART models, almost always outperform their respective linear counterparts. Finally, the results from Tab. 4 and 5 are confirmed even in the presence of partial ranking.